# Lattice points on determinant surfaces and the spectrum of the automorphic Laplacian

### v1

@@ 7.3. Subdividing the range of integration and decomposition of the sum $\mathcal{S}$
Note that for $11\leq j\leq k+$ 1, the interval $\mathcal{I}_{j}$ is contained in $\mathcal{F}_{j}$ and the edges of $\mathcal{I}_{j}$ are away from the edges of $\mathcal{F}_{j}$. For isolating the stationary phase when the stationary point $x_{0}\in\mathcal{I}_{j}$ we express the sum $\mathcal{S}$ as
@@
Here the main idea is to subdivide the ranges of $m$, $n$ and $a_{1}$ according to the position of $x_{0}$. First we interchange the sum over $j$ and the sum over $a_{1}$ using (7.4) and (7.9). Then we bring the sum over $j$ outside the sums over the variables $m$ and $n$. Thus we have,
@@
Note that the sum over $a_{1}$ is vacuous if $j$ is so large that $X/2^{j}<la_{1}$. We now make a dyadic subdivision of the range of $x_{0}$ and write
$[y_{j}+\delta,X^{3/2}]=\bigcup_{\begin{subarray}{c}Y=2^{j}\delta\\ 0\leq j\leq L(X)\end{subarray}}[y_{j}+Y,y_{j}+2Y],$
@@
$\frac{n}{m}\leq\frac{a_{1}X}{(y_{j}+Y)^{2}}\ll\frac{1}{l}\left(\frac{X}{y_{j}}\right)^{2},$ (7.14)
and
$\frac{n}{mX}(y_{j}+Y)^{2}\leq a_{1}\leq\frac{n}{mX}(y_{j}+2Y)^{2},$ (7.15)
@@
Case II: $Y>\frac{y_{j}}{2}$. By (7.1), the first inequality of (7.16) and Prop. 5.5, this case contributes $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ in $Q_{1}(X)$ (see (7.12)). Combining this with the bound in Case I, we prove Prop. 7.3. ∎
@@ 7.5. Estimation of $\mathcal{T}$
say. By an argument similar to the one used in proving Prop. 7.3, we can show that
@@
Since $x_{0}\in\mathcal{I}_{j}$, and $\mathcal{I}_{j}\subset\mathcal{F}_{j}$, we extract the stationary phases by applying Lemma 2.6 and thus obtain that for $11\leq j\leq k+$ 1,
@@
and everything else remains unchanged. As in the proof of Prop. 7.3, by introducing dyadic subdivision of the range of the stationary point and analyzing the ensuing restrictions on the variables $m,n$ and $a_{1}$, we show that the contributions of the error terms in $\mathcal{S}_{3}$, $\mathcal{S}_{5}$ and $\mathcal{S}_{7}$ is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the Weil bound. Furthermore, from (7.19), we observe that the stationary phase analysis of the integrals in $\mathcal{S}_{3},\mathcal{S}_{5},\mathcal{S}_{7}$ give the same main term. Combining all these estimates in one place, we conclude that
$Q_{3}(X):=\mathcal{S}_{3}+\mathcal{S}_{5}+\mathcal{S}_{7}$
@@
Now we extend the range of $x_{0}$ to the whole interval $[la_{1},X]$ and subtract the contribution of the extra terms that have added. Note that the total contribution of these extra terms is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the same kind of argument used in the proof of Prop. 7.4. Thus, we have
@@
Combining Prop. 7.3, Prop. 7.4, (7.20) and also considering the case $m,n<0$, we prove Prop. 7.1.
@@ 7.7. Contribution of $I_{5}$
By Prop. 5.5, the contribution of the $O$-term in $B_{1,w}(X)$ is $O\left(X^{3/2+\varepsilon}\right)$.
@@
Bringing together Prop. 6.1, Prop. 7.1 and Prop. 7.5, we arrive at Prop. 5.6, except for the conditions on $a_{1}$. Now, the condition $la_{1}\leq\sqrt{\frac{ma_{1}X}{n}}\leq X$ appearing in the sum over $a_{1}$ in Prop. 7.1 is equivalent to saying $0<a_{1}\leq\min\{\frac{mX}{nl^{2}},\frac{nX}{m},\frac{X}{l}\}$. However, we can drop the condition $a_{1}\leq\frac{X}{l}$ because of the presence of the factor $w(la_{1})$ in our sum and in view of Remark 5.4. A similar consideration shows that the condition $la_{1}\leq\sqrt{\frac{ma_{1}X}{n}}\leq X$ is equivalent to saying $0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}$.
Remark 7.6.
At this stage, a direct application of the Weil bound for Kloosterman sum (see Lemma 2.3), yields the bound
Now, from (3.2), (3.9), (3.10), and Prop. 3.2, the optimal choice for $H$ is given by $H=X^{2/3}$ and this gives us
${S_{r}}(X)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(X^{5/3+\varepsilon}\right).$
8. Removal of the sharp-cut condition and the proof of Proposition 3.3
To simplify notation, let us, henceforth, denote the variable $a_{1}$ by $c$.
In Prop. 5.6, we expressed $\widehat{G}(n/c,m/c)$ in a form that enabled us to see its oscillating behaviour and its order of magnitude. The next step is to apply the Kuznetsov formula to the sums over the variable $c$, but first we need to make the variable smooth. We show how to do it for $T_{1,w}(X)$. The condition $c\leq\min\left\{\frac{mX}{nl^{2}},\frac{nX}{m}\right\}$ can be restated as
@@
$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{0<c\leq\frac{nX}{m}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{mc}{n}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
@@
$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m\leq nl}\sum_{0<c\leq\frac{mX}{nl^{2}}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m\leq nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{cnl^{2}}{m}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
@@
$R_{1,w}(X)$ $:=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right),$
@@
${R^{\prime}}_{1,w}(X)$ $:=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{m<nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{nl^{2}c}{m}\right).$
@@
$R_{3,w}(X)$ $:=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right),$ (8.4)
@@
${R^{\prime}}_{3,w}(X)$ $:=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l}\sum_{m}\sum_{n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{nl^{2}c}{m}\frac{X}{u}\right).$
The sum $T_{2,w}(X)$ (resp. $T_{4,w}(X)$) is exactly analogous to $T_{1,w}(X)$ (resp. $T_{3,w}(X)$) and we have similar decompositions for $T_{2,w}(X)$ and $T_{4,w}(X)$. Combining these expressions for $T_{i,w}(X),i=1,2,3,4$, we obtain Prop. 3.3.
@@
We only discuss the cases of $T_{1,w}(X)$ and $T_{3,w}(X)$ below as the other two sums can be handled similarly. Moreover, it is enough to consider $R_{1,w}(X)$ and $R_{3,w}(X)$ as ${R^{\prime}}_{1,w}(X)$ and ${R^{\prime}}_{3,w}(X)$ are similar to $R_{1,w}(X)$ and $R_{3,w}(X)$, respectively.
@@ 9.1. Analysis of $R_{1,w}(X)$
$R_{1,w}(X)$ $=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}e\left(-\frac{1}{8}\right)\sum_{l|r}\mathop{\sum_{M\text{ dyadic}}\sum_{N\text{ dyadic}}\sum_{C\text{ dyadic}}}\mathop{\sum_{1\leq m\ll\frac{X^{1+\varepsilon}}{H}}\sum_{1\leq n\ll\frac{X^{1+\varepsilon}}{lH}}}_{nl\leq m}b_{1}\left(\frac{m}{M}\right)b_{2}\left(\frac{n}{N}\right)$ (9.1)
×$\times\sum_{c}V\left(\frac{c}{C}\right)\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right).$
When $C$ is small enough, the Weil bound already gives a good bound (see Prop. 5.5) and we collect all such terms in the error term and write
$R_{1,w}(X)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}e\left(-\frac{1}{8}\right)\sum_{l|r}\mathop{\sum_{M\text{ dyadic}}\sum_{N\text{ dyadic}}\sum_{\begin{subarray}{c}C\text{ dyadic}\\ C\geq\frac{X^{\frac{4}{5}+\varepsilon}}{l}\end{subarray}}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots+O\left(X^{\frac{5}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{1}{2}}\right).$ (9.2)
Now we fix a divisor $l$ of $r$ and consider the sum
$R_{1}(M,N,C,X):=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right)V\left(\frac{c}{C}\right)g(m,n,c),$
@@
$g(m,n,c):=\left(\frac{c}{mn}\right)^{\frac{3}{4}}.$ (9.3)
We define
@@
so that the contribution of the terms corresponding to $C\leq C_{0}$ is absorbed in the error term after we make the choice $H=\sqrt{X}$ at the end. Note that in the sum $R_{1}(M,N,C,X)$, the dyadic variables $M$ and $N$ satisfy $1\leq M\ll\frac{X^{1+\varepsilon}}{H}$, $1\leq N\ll\frac{X^{1+\varepsilon}}{lH}$ and from the support of the function $w$ (see (3.6)), we have $C\leq\frac{8NX}{M}$ as $N\leq n\leq 4N$ and $M\leq m\leq 4M$. For estimating $R_{1,w}(X)$, it is enough to estimate $R_{1}(M,N,C,X)$ and here is the main result we need.

### v2

@@ 7.3. Subdivision of the range of integration and decomposition of the sum $\mathcal{S}$
Note that for $1\leq j\leq k+1$, the interval $\mathcal{I}_{j}$ is contained in $\mathcal{F}_{j}$ and the edges of $\mathcal{I}_{j}$ are away from the edges of $\mathcal{F}_{j}$. For isolating the stationary phase when the stationary point $x_{0}\in\mathcal{I}_{j}$ we express the sum $\mathcal{S}$ as
@@
Here the main idea is to subdivide the ranges of $m$, $n$ and $a_{1}$ according to the position of $x_{0}$. Interchanging the order of summations and using (7.4) and (7.9), we have
@@
Note that the sum over $a_{1}$ is vacuous if $j$ is so large that $X/2^{j}<la_{1}$. Now write
$y_{j}+\delta,X^{3/2}]=\bigcup_{\begin{subarray}{c}Y=2^{j}\delta\\ 0\leq j\leq L(X)\end{subarray}}[y_{j}+Y,y_{j}+2Y],$
@@
$\frac{n}{m}\leq\frac{a_{1}X}{(y_{j}+Y)^{2}}\ll\frac{1}{l}\left(\frac{X}{y_{j}}\right)^{2},\quad\textnormal{ and }\quad\frac{n}{mX}(y_{j}+Y)^{2}\leq a_{1}\leq\frac{n}{mX}(y_{j}+2Y)^{2},$ (7.14)
@@
Case II: $Y>\frac{y_{j}}{2}$. By (7.1), the first inequality of (7.15) and Prop. 5.4, this case contributes $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ in $Q_{1}(X)$ (see (7.12)). Combining this with the bound in Case I, we prove Prop. 7.2. ∎
@@ 7.5. Estimation of $\mathcal{T}$
say. By an argument similar to the one used in proving Prop. 7.2, it follows that
@@
Since $x_{0}\in\mathcal{I}_{j}$, and $\mathcal{I}_{j}\subset\mathcal{F}_{j}$, we extract the stationary phases by applying Lemma 2.9 and thus obtain that for $1\leq j\leq k+1$,
@@
and everything else remains unchanged. As in the proof of Prop. 7.2, by introducing dyadic subdivision of the range of the stationary point and analyzing the ensuing restrictions on the variables $m,n$ and $a_{1}$, we show that the contributions of the error terms in $\mathcal{S}_{3}$, $\mathcal{S}_{5}$ and $\mathcal{S}_{7}$ is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the Weil bound. Furthermore, from (7.18), we observe that the stationary phase analysis of the integrals in $\mathcal{S}_{3},\mathcal{S}_{5},\mathcal{S}_{7}$ give the same main term. Combining all these estimates in one place, we conclude that
$Q_{3}(X)=\mathcal{S}_{3}+\mathcal{S}_{5}+\mathcal{S}_{7}$
@@
Now we extend the range of $x_{0}$ to the whole interval $la_{1},X]$ and subtract the contribution of the extra terms that have been added. Note that the total contribution of these extra terms is $O\left(X^{\frac{3}{2}+\varepsilon}\right)$ by the same kind of argument used in the proof of Prop. 7.3. Thus, we have
@@
Combining Prop. 7.2, Prop. 7.3, (7.19) and also considering the case $m,n<0$, we prove Prop. 7.1.
@@ 7.7. Contribution of $I_{5}$
By Prop. 5.4, the contribution of the $O$-term in $B_{1,w}(X)$ is $O\left(X^{3/2+\varepsilon}\right)$.
@@
Bringing together Prop. 6.1, Prop. 7.1 and Prop. 7.4, we arrive at Prop. 5.1, except for the conditions on $a_{1}$. Now, the condition $la_{1}\leq\sqrt{\frac{ma_{1}X}{n}}\leq X$ appearing in the sum over $a_{1}$ in Prop. 7.1 is equivalent to saying $0<a_{1}\leq\min\{\frac{mX}{nl^{2}},\frac{nX}{m},\frac{X}{l}\}$. However, we can drop the condition $a_{1}\leq\frac{X}{l}$ because of the presence of the factor $w(la_{1})$ in our sum and in view of Remark 5.1. A similar consideration shows that the condition $la_{1}\leq\sqrt{\frac{ma_{1}X}{n}}\leq X$ is equivalent to saying $0<a_{1}\leq\min\{\frac{mu}{nl^{2}},\frac{nX^{2}}{mu}\}$.
7.9. Removal of the sharp-cut condition and the proof of Proposition 3.2
To simplify notation, let us, henceforth, denote the variable $a_{1}$ by $c$ with the understanding that this has no connection the original variable $c$.
In Prop. 5.1, we expressed $\widehat{G}(n/c,m/c)$ in a form that enabled us to see its oscillating behaviour and its order of magnitude. We now need to make the variable smooth and we show how to do it for $T_{1,w}(X)$. The condition $c\leq\min\left\{\frac{mX}{nl^{2}},\frac{nX}{m}\right\}$ can be restated as
@@
$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{nl\leq m}\sum_{0<c\leq\frac{nX}{m}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{mc}{n}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
@@
$\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{m\leq nl}\sum_{0<c\leq\frac{mX}{nl^{2}}}\frac{S(nr_{1},-m,c)}{c}w(lc)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)$
$=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{m\leq nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}w\left(\frac{cnl^{2}}{m}\right)\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)+O\left(X^{1+\varepsilon}\sqrt{H}\right).$
@@
$R_{1,w}(X)$ $=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{nl\leq m}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right),$
@@
${R^{\prime}}_{1,w}(X)$ $=\frac{X^{\frac{1}{4}}e(-1/8)}{2\sqrt{2}\pi i}\mathop{\sum_{l,m,n}}_{m<nl}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{3}{4}}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{nl^{2}c}{m}\right).$
@@
$R_{3,w}(X)$ $=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l,m,n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{mc}{n}\frac{u}{X}\right),$ (7.24)
@@
${R^{\prime}}_{3,w}(X)$ $=\frac{e(-1/8)}{\sqrt{2}}\int\limits_{X}^{X+H}\frac{w(u)}{u^{1/4}}\mathop{\sum_{l,m,n}}_{nl\leq m\frac{u}{X}}\sum_{c>0}\frac{S(nr_{1},-m,c)}{c}\left(\frac{c}{mn}\right)^{\frac{1}{4}}e\left(-2\sqrt{\frac{mnu}{c}}\right)w\left(\frac{nl^{2}c}{m}\frac{X}{u}\right).$
The sum $T_{2,w}(X)$ (resp. $T_{4,w}(X)$) is exactly analogous to $T_{1,w}(X)$ (resp. $T_{3,w}(X)$) and we have similar decompositions for $T_{2,w}(X)$ and $T_{4,w}(X)$. Combining these expressions for $T_{i,w}(X),i=1,2,3,4$, we obtain Prop. 3.2.
@@
It is enough to consider the sums $T_{1,w}(X)$ and $T_{3,w}(X)$ as the other two sums can be handled similarly. Moreover, it is enough to consider $R_{1,w}(X)$ and $R_{3,w}(X)$ as ${R^{\prime}}_{1,w}(X)$ and ${R^{\prime}}_{3,w}(X)$ are similar to the other two sums.
@@ 8.1. Analysis of $R_{1,w}(X)$
$R_{1,w}(X)=\frac{X^{\frac{1}{4}}}{2\sqrt{2}\pi i}e\left(-\frac{1}{8}\right)\sum_{l|r}\mathop{\sum_{M\text{ dyadic}}\sum_{N\text{ dyadic}}\sum_{\begin{subarray}{c}C\text{ dyadic}\\ C\geq\frac{X^{\frac{4}{5}+\varepsilon}}{l}\end{subarray}}}\mathop{\sum_{m}\sum_{n}}_{nl\leq m}\cdots+O\left(X^{\frac{5}{4}+\varepsilon}\left(\frac{X}{H}\right)^{\frac{1}{2}}\right),$ (8.1)
where we have collected the contributions of the smaller $c$’s in the error term after applying the Weil bound (see Prop. 5.4). It is, therefore, enough to estimate the sum
$R_{1}(M,N,C,X)=\mathop{\sum_{m}\sum_{n}}_{nl\leq m}b_{1}(m/M)b_{2}(n/N)\sum_{c}\frac{S(nr_{1},-m,c)}{c}e\left(-2\sqrt{\frac{mnX}{c}}\right)w\left(\frac{mc}{n}\right)V\left(\frac{c}{C}\right)g(m,n,c),$
@@
$g(m,n,c):=\left(\frac{c}{mn}\right)^{\frac{3}{4}}$ (8.2)
and $l$ is a fixed divisor of $r$. Let us define
@@
so that the contribution of the terms corresponding to $C\leq C_{0}$ is absorbed in the error term after we make the choice $H=\sqrt{X}$ at the end. Recall that $1\leq M\ll\frac{X^{1+\varepsilon}}{H}$, $1\leq N\ll\frac{X^{1+\varepsilon}}{lH}$ and from (3.6), we have $C\leq\frac{8NX}{M}$ since $N\leq n\leq 4N$ and $M\leq m\leq 4M$.